OpenStudy (anonymous):

solve using substitution 9x+9y=-27 -5x+y=39

OpenStudy (anonymous):

then use substitution.. those are just simple equations..

OpenStudy (anonymous):

divide first equation by 3 to start - it will make things easier

OpenStudy (anonymous):

i would change the 2nd equation to y=5x+39

OpenStudy (anonymous):

i mean 9 !!!!

OpenStudy (anonymous):

x+y=-3

OpenStudy (anonymous):

right jimmy?

OpenStudy (anonymous):

it then becomes x + y = -3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y=-3-x sub in other one -5x-(-3-x)=39 -5x+3+x=39 -4x=36 x=-9

OpenStudy (anonymous):

now plug into x=-9 into y=-3-x

OpenStudy (anonymous):

you get y=6

OpenStudy (anonymous):

thanks :)